the point diametrically opposite to the point p(1,0) on the circle x2+y2+2x+3y

Grade:11

4 Answers

Anoopam Mishra
8 years ago
Let the point be (h,k), then the centre of the circle is ((h+1)/2,k/2)
But the centre of the circle is (-2,-3) from the givern equation.
=> (h+1)/2 = -2 => h=-5
and k/2 = -3 => k=-6.
Hence the diametrically opposite point is (-5,-6)
Raunaq Mehta
7 years ago
Let the point be (h,k), then the centre of the circle is ((h+1)/2,k/2)But the centre of the circle is (-2,-3) from the givern equation.=> (h+1)/2 = -2 => h=-5and k/2 = -3 => k=-6.Hence the diametrically opposite point is (-5,-6)
Shubh mehta
6 years ago
Let the point be (h,k), then the centre of the circle is ((h+1)/2,k/2)But the centre of the circle is (-2,-3) from the givern equation.=> (h+1)/2 = -2 => h=-5and k/2 = -3 => k=-6.so the diametrically opposite point is (-5,-6)
Abrar gani
6 years ago
X2+y2 +2x+4y-3=oC(-1,-2)As P and Q are diametrically opposite therefore point C is the midpoint of line joining PandeyLetQ(x, y)x+1/2=1x=-3andy+0/2=-2Y=-4Implies that Point Q(-3,-4)